OpenStudy (anonymous):

I need to find an angle in the third quadrant where sin(theta) = -1/2 and cos(theta) = 1/2

OpenStudy (anonymous):

There is no such angle. \((\frac{1}{2})^2 + (-\frac{1}{2})^2 \not = 1^2\).